The radius and surface charge density of a uniformly charged spherical shell are

20 cm and ,m C0.3 −2  respectively. Calculate the electric potential at a distance

(a) 40 cm and (b) 15 cm from the center of the shell.


1
Expert's answer
2021-06-21T16:02:46-0400
"q=4\pi(0.2)^2(0.3\cdot10^{-6})=1.5\cdot10^{-7}C"

a)


"V=9\cdot10^{9}\frac{1.5\cdot10^{-7}}{0.4}=3375\ V"

b)


"V=9\cdot10^{9}\frac{1.5\cdot10^{-7}}{0.2}=6750\ V"